OpenStudy (anonymous):

In a club, 3 members bring all of the sandwitches for a picnic. Mike and lynn arive together. They have allready put together mike's 3 sandwithces and Lynn's 5 sandwitches. Hosea comes a little later with 4 sandwitches. In all, the club has a total of 12 sandwitches. write an addition expression that shows the grouping described above. Then suppose

OpenStudy (anonymous):

*already *sandwiches

OpenStudy (anonymous):

That mike had come first, alone, and that Lynn and hosea had come with their sandwitches later. Write an addition expression to represent this grouping.

OpenStudy (anonymous):

Also Would the club's total number of sandwitches be the same with either grouping?

OpenStudy (anonymous):

(3+5)+4=12

OpenStudy (anonymous):

so this is an example of associativity. If they arrive together, the number of sandwiches arriving with the people who arrived together would be in brackets

OpenStudy (anonymous):

the answer would always be twleve

OpenStudy (anonymous):

yup, always twelve. that's what associativity means

OpenStudy (anonymous):

(3+5)+4=12 thats right tho right?

OpenStudy (anonymous):

right :-p

OpenStudy (anonymous):

well, depending on who arrives together. I've not really read the nitty gritty bits of the question

OpenStudy (anonymous):

Write an addition expression that shows the grouping described above. And write an additon expression to represent this grouping. ?

OpenStudy (anonymous):

so if the people arrive together, put their sandwiches in brackets (you know what i mean)

OpenStudy (anonymous):

ok so ( 3+5) +4 = 12

OpenStudy (anonymous):

what abt the second part of the question?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just reorder the brackets

OpenStudy (anonymous):

(5+4)+3=12